### v1

@@ 4.2 Camera Sampling
In the first stage, we generate a coarse representation of the surrounding environments (indoor walls and outdoor faraway surroundings). We freeze the 3D Gaussians parameters of the ground and objects, sampling camera coordinates within a certain range of the scene center to optimize the generation of the surrounding environments.
In the second stage, we focus on generating the coarse ground. We freeze the 3D Gaussians parameters of the environments and objects. The indoor scene is divided into different regions based on the placement of objects, we sample camera poses pointing at these focal regions containing objects and the ground of the room in each iteration. The outdoor scene is divided into concentric circles based on radius, and in each iteration, a same direction is chosen to sample camera poses on different circles to optimize ground generation. To enhance the uniformity and consistency between the ground and its surrounding environments, our strategy is able to cover the entire ground as uniformly as possible and focuses on optimizing the parts where the ground is in contact with objects and the surrounding environments.
@@ 5 Experiment
Implementation Details. We utilize GPT-4 as our LLM for scene prompt decomposition, Point-E [24] generated sparse point clouds for the initial representation of objects, Stable Diffusion 2.1 as our 2D text-to-image model. The maximum number of iterations is set to 1,500 and 2,000 rounds for objects and the environment, respectively. The initial time interval value $m$, starts at 4 and decreases by 1 every 400 rounds. The number of rendering images in the reconstructive generation is set to 20. We tested DreamScene and all the baselines on the same NVIDIA 3090 GPU for fair comparison. Baselines. For the comparison of the text-to-3D scene generation, we use the current open-sourced SOTA methods Text2Room [9], Text2NeRF [40], ProlificDreamer [37], and Set-the-Scene [3] as the baselines. For the comparison of typical text-to-3D generation, we select the current open-source SOTA methods DreamFusion [28], Magic3D [17], DreamGaussian [35], and LucidDreamer [16] as the baselines (ProlificDreamer[37], DreamFusion [28] and Magic3D [17] were reimplemented by Three-studio [5]). Evaluation Metrics. We tested the generation time of each method [9, 40, 3, 37, 28, 17, 35, 16], compared the editing capabilities against their published papers, and did a 100-participant user study to score (out of 5) the quality, consistency, and rationality of the videos generated by each method for 5 scenes (3 indoor, 2 outdoor) of 30 seconds.
@@ 5.1 Qualitative Results
Fig. 4(a) shows the comparison of DreamScene in indoor and outdoor scenes with the current SOTA methods[11, 40, 37, 3]. The upper images are rendered with the camera poses that appeared during the generation, and the lower images are rendered with the randomly selected camera poses within the scene. We can see that Text2Room[9] and Text2NeRF[40] only produce satisfactory results under camera poses encountered during generation. Combined with the results from Fig. 5, it shows that DreamScene achieves the best 3D consistency alongside commendable generation quality. For the comparison of generating single objects with text-to-3D methods, Fig. 4(b) shows that our FPS can generate high-quality 3D representation following the text prompts in a short time. DreamGaussian [35] is faster yet at the cost of too low generation quality. Please refer to the Supplementary Materials for additional images and video results of scene and object generation.
@@ 5.2 Quantitative Results
Figure 8: The ablation results of the 3D Gaussian filtering algorithm in reconstruction and generation tasks.
@@ 5.4 Ablations
We compared the effects of different sampling strategies via the generation results of a 3D object. Fig. 7 shows the results optimized for 30 minutes under the prompt “A DSLR photo of Iron Man”. As illustrated, multi-timestep sampling (MTS) forms better geometric structures and textures compared to the monotonically non-increasing sampling in DreamTime [10] and the Score Distillation Sampling (SDS) mentioned in DreamFusion [28]. Formation Pattern Sampling (FPS), building on the foundation of MTS, employs a reconstruction method to create smoother and more plausible textures that demonstrate the superiority of FPS. The 3D Gaussian filtering we proposed is specifically designed for optimization tasks and can also be directly applied to reconstruction tasks [13]. Fig. 8 compares the results of reconstruction and generation tasks before and after compression using the Gaussian filtering algorithm. It can be seen that in the reconstruction task, our compression achieved 73.9%, with the overall image slightly blurred and some details lost. In our generation task, however, the compression was 66.1% without significant quality loss. Due to space constraints, we have placed the ablation experiments on camera sampling and more comprehensive results in the Supplementary Materials.
@@ 6 Applications, Limitation and Conclusion
DreamScene has the potential to impact numerous industries, as illustrated in Fig. 9. For instance, in VR gaming and the Metaverse, DreamScene, integrated with LLMs, enables the capability to generate a fantasy scene merely from a user’s description. Moreover, it allows for easy modification of object positions and the style of the scene through descriptive means. DreamScene can also revolutionize film production by enabling rapid and customizable generation of intricate 3D scenes directly from scripts, significantly enhancing visual storytelling and reducing the reliance on physical sets and manual modeling. For architectural design, DreamScene facilitates the generation of furniture in various styles, enabling users to flexibly design and arrange objects within a scene using layout diagrams.
@@ 8.1 Rendering and Training Settings
19:              $\epsilon_{\phi}(x_{i};y_{n},t_{i})=$U-Net$(x_{i},y,t_{i})$
20:              $\epsilon_{\phi}(x_{i};\emptyset,t_{i})=$U-Net$(x_{i},\emptyset,t_{i})$
@@
22:         $\nabla_{\theta}\mathcal{L}_{\text{MTS}}(\theta)=\mathbb{E}_{t,\epsilon,c}\left[\sum\limits_{i=1}^{m}w(t)(\epsilon_{\phi}(x_{t_{i}};y_{n},t_{i})-\epsilon_{\phi}(x_{t_{i}};\emptyset,t_{i}))\frac{\partial g(\theta,c)}{\partial\theta}\right]$
@@
24:         if $iter\%compress\_iter$ then

### v2

@@ 4.2 Camera Sampling
In the first stage, we generate a coarse representation of the surroundings (indoor walls and outdoor faraway surroundings). We freeze the 3D Gaussians parameters of the ground and objects, sampling camera coordinates within a certain range of the center to optimize the generation of the surroundings.
The second stage focuses on generating the coarse ground. We freeze the 3D Gaussians parameters of the environments and objects. The indoor scene is divided into regions based on the placement of objects, we sample camera poses pointing at these focal regions containing objects and the ground of the room in each iteration. The outdoor scene is divided into concentric circles based on radius, and in each iteration, a same direction is chosen to sample camera poses on different circles to optimize ground generation. Our strategy is able to cover the entire ground as uniformly as possible and focuses on optimizing the parts where the ground is in contact with objects and the surroundings.
@@ 5 Experiment
Implementation Details. We utilize GPT-4 as our LLM for scene prompt decomposition, Point-E [25] generated sparse point clouds for the initial representation of objects, Stable Diffusion 2.1 as our 2D text-to-image model. The maximum number of iterations is set to 1,500 and 2,000 rounds for objects and the environment, respectively. The initial time interval value $m$, starts at 4 and decreases by 1 every 400 rounds. The number of rendering images in the reconstructive generation is set to 20. We tested DreamScene and all the baselines on the same NVIDIA 3090 GPU for fair comparison. Baselines. For the comparison of the text-to-3D scene generation, we use the current open-sourced SOTA methods Text2Room [10], Text2NeRF [41], ProlificDreamer [38], and Set-the-Scene [4] as the baselines. For text-to-3D generation, we select open-source SOTA methods DreamFusion [29], Magic3D [18], DreamGaussian [36], and LucidDreamer [17] as the baselines (ProlificDreamer, DreamFusion and Magic3D were reimplemented by Three-studio [6]). Evaluation Metrics. We tested the generation time of each method [10, 41, 4, 38, 29, 18, 36, 17], compared the editing capabilities against their published papers, and did a 100-participant user study to score (out of 5) the quality, consistency, and rationality of the videos generated by each method for 5 scenes (3 indoor, 2 outdoor) of 30 seconds.
@@ 5.1 Qualitative Results
Fig. 4(a) shows the comparison of DreamScene in indoor and outdoor scenes with the SOTA methods[12, 41, 38, 4]. The upper images are rendered with the camera poses that appeared during the generation, and the lower ones with the randomly selected camera poses within the scene. We can see that Text2Room[10] and Text2NeRF[41] only produce satisfactory results under camera poses encountered during generation. Combined with the results from Fig. 5, it shows that DreamScene achieves the best 3D consistency alongside commendable generation quality. For the comparison of generating single objects with text-to-3D methods, Fig. 4(b) shows that our FPS can generate high-quality 3D representation following the text prompts in a short time. DreamGaussian [36] is faster yet at the cost of too low generation quality. Please refer to the Supplementary Materials for additional images and video results.
@@ 5.2 Quantitative Results
Figure 8: Ablation results of 3D Gaussian filtering algorithm in reconstruction and generation tasks.
@@ 5.4 Ablations
We compared the effects of different sampling strategies via the generation results of a 3D object. Fig. 7 shows the results optimized for 30 minutes under the prompt “A DSLR photo of Iron Man”. As illustrated, multi-timestep sampling (MTS) forms better geometric structures and textures compared to the monotonically non-increasing sampling in [11] and the SDS in [29]. Formation Pattern Sampling (FPS), building on top of MTS, employs a reconstruction method to create smoother and more plausible textures. The 3D Gaussian filtering we proposed is specifically designed for optimization tasks and can also be directly applied to reconstruction tasks [14]. Fig. 8 compares the results of reconstruction and generation before and after compression using the Gaussian filtering algorithm. It can be seen that in the reconstruction task, our compression achieved 73.9%, with the overall image slightly blurred and some details lost. In our generation task, however, the compression was 66.1% without significant quality loss. Please find the ablation experiments on camera sampling and more comprehensive results in the Supplementary Materials.
@@ 6 Applications, Limitation and Conclusion
DreamScene can impact numerous industries as illustrated in Fig. 9. For instance, in VR gaming and the Metaverse, DreamScene, integrated with LLMs, enables the capability to generate a fantasy scene merely from a user’s description. Moreover, it allows for easy modification of object positions and the style of the scene through descriptive means. DreamScene can also revolutionize film production by enabling rapid and customizable generation of intricate 3D scenes directly from scripts, significantly enhancing visual storytelling and reducing the reliance on physical sets and manual modeling. For architectural design, DreamScene facilitates the generation of furniture in various styles, enabling users to flexibly design and arrange objects within a scene using layout diagrams.
@@
Acknowledgements: This work was supported by the National Key Research and Development Program of China (2022YFB3105405, 2021YFC3300502).
@@ 8.1 Rendering and Training Settings
19:              $\epsilon_{\phi}(x_{t_{i}};y_{n},t_{i})=$U-Net$(x_{t_{i}},y_{n},t_{i})$
20:              $\epsilon_{\phi}(x_{t_{i}};\emptyset,t_{i})=$U-Net$(x_{t_{i}},\emptyset,t_{i})$
@@
22:         $\nabla_{\theta}\mathcal{L}_{\text{MTS}}(\theta)=\mathbb{E}_{t,\epsilon,c}\left[\sum\limits_{i=1}^{m}w(t_{i})(\epsilon_{\phi}(x_{t_{i}};y_{n},t_{i})-\epsilon_{\phi}(x_{t_{i}};\emptyset,t_{i}))\frac{\partial g(\theta,c)}{\partial\theta}\right]$
@@
24:         if $iter\%compress\_iter=0$ then